# biblio.bib

### v2

@@ 11 Six points
$X_{1:5}=\begin{pmatrix}1&A&A&A&A&A\\ A&1&B&C&C&B\\ A&B&1&B&C&C\\ A&C&B&1&B&C\\ A&C&C&B&1&B\\ A&B&C&C&B&1\end{pmatrix},\quad\begin{array}[]{l}A:=-\frac{1}{5},\\ B:=\frac{-5+6\sqrt{5}}{25},\\ C:=\frac{-5-6\sqrt{5}}{25}\end{array}.$
$X_{1:4:1}=\begin{pmatrix}1&A&0&0&0&0\\ A&1&0&0&0&0\\ 0&0&1&0&A&0\\ 0&0&0&1&0&A\\ 0&0&A&0&1&0\\ 0&0&0&A&0&1\end{pmatrix},\quad A:=-1.$
$X_{\text{Eq.}}=\begin{pmatrix}1&A&B&C&B&A\\ A&1&A&B&C&B\\ B&A&1&A&B&C\\ C&B&A&1&A&B\\ B&C&B&A&1&A\\ A&B&C&B&A&1\end{pmatrix},\quad\begin{array}[]{l}A:=\frac{1}{2},\\ B:=-\frac{1}{2},\\ C:=-1\end{array}$
@@
11.1 Configuration 3:3 and its real conjugate
@@
$X_{3:3}=\begin{pmatrix}1&A&A&B&C&C\\ A&1&A&C&B&C\\ A&A&1&C&C&B\\ B&C&C&1&A&A\\ C&B&C&A&1&A\\ C&C&B&A&A&1\end{pmatrix},\quad\begin{array}[]{l}A:=\frac{-7+3\sqrt{6}}{5}\\ B:=\frac{11-4\sqrt{6}}{5}\\ C:=\frac{-1-\sqrt{6}}{5}\end{array}.$
To verify that $X_{3:3}$ is critical, we replace its values in the polynomial equations using M2. As the matrix depends on $\sqrt{6}$, we define a polynomial ring with coefficients in $k:=\mathbb{Q}[t]/(t^{2}-6)$. This introduces a variable $t$, such that $t^{2}=6$. We then replace $\sqrt{6}$ by $t$ in $X_{3:3}$, and check that its entries verify the equations in $k[x_{ij},z_{ij}]$. As this is true for every $t$ such that $t^{2}=6$, it implies that if we change $\sqrt{6}$ by $-\sqrt{6}$ in $X_{3:3}$, we also obtain a solution to the polynomial equations. We call this the “real conjugate” of 3:3, with matrix:
$X_{\overline{3:3}}=\begin{pmatrix}1&\overline{A}&\overline{A}&\overline{B}&\overline{C}&\overline{C}\\ \overline{A}&1&\overline{A}&\overline{C}&\overline{B}&\overline{C}\\ \overline{A}&\overline{A}&1&\overline{C}&\overline{C}&\overline{B}\\ \overline{B}&\overline{C}&\overline{C}&1&\overline{A}&\overline{A}\\ \overline{C}&\overline{B}&\overline{C}&\overline{A}&1&\overline{A}\\ \overline{C}&\overline{C}&\overline{B}&\overline{A}&\overline{A}&1\end{pmatrix},\quad\begin{array}[]{l}\overline{A}:=\frac{-7-3\sqrt{6}}{5}\simeq-2.87\\ \overline{B}:=\frac{11+4\sqrt{6}}{5}\simeq 4.16\\ \overline{C}:=\frac{-1+\sqrt{6}}{5}\simeq 0.29\end{array}.$
As $X_{\overline{3:3}}$ has entries with absolute value greater than one, it is not the dot product of points on a unit sphere (it corresponds to a complex critical configuration). Note also that we already encountered configurations with non-rational values: the Equator for $n=5$, and 1:5; both depending on $\sqrt{5}$. However, in those cases, changing $\sqrt{5}$ by $-\sqrt{5}$ gives a matrix in the same orbit as the original.
@@
11.2 Possible values for the dot product
@@ 8 $(5x_{45}^{2}+1)^{2}$$\pm\frac{i}{\sqrt{5}}$
9 $55x_{45}^{2}-22x_{45}+$ 5 $\frac{11\pm 4\sqrt{6}}{5}$
10 $15x_{45}^{2}+2x_{45}-$ 1 $\frac{-1\pm\sqrt{6}}{5}$
11 $15x_{45}^{2}+14x_{45}-$ 1 $\frac{-7\pm 3\sqrt{6}}{5}$
12 $1925x_{45}^{2}+28x_{45}+$ 19 $\frac{-14\pm i3\sqrt{31}}{25}$
13 $31125x_{45}^{2}+50x_{45}-$ 31 $\frac{-5\pm 6\sqrt{5}}{25}$
14 $10100x_{45}^{4}+95x_{45}^{3}-21x_{45}^{2}-22x_{45}+$ 10 4 complex
15 $4250x_{45}^{4}+110x_{45}^{3}-21x_{45}^{2}-19x_{45}+$ 4 4 complex
16 $67400x_{45}^{4}+488x_{45}^{3}-111x_{45}^{2}-196x_{45}+$ 67 4 complex
17 $13x_{45}+$ 1 -1/3
18 $45x_{45}+$ 4 -4/5
19 $110x_{45}-$ 1 1/10
20 $125x_{45}-$ 1 1/25
21 $1125x_{45}+$ 11 -11/25
22 $2325x_{45}+$ 23 -23/25
11.3 Minimal primes and new solutions
@@
11.4 Orbits including new solutions
@@
Difference $938-848=90$
11.5 Multiplicity of “Complex 1”
Proposition 4 ([hartshorne2013algebraic, Theorem 5.1]).
Consider a polynomial ideal $I\subset\mathbb{Q}[x_{1},\ldots,x_{m}]$, generated by polynomials $\{g_{1},\ldots,g_{r}\}$. Define the Jacobian matrix of the ideal generators as:
$J(x)\in\mathbb{R}^{r\times m}\quad/\quad J(x)_{ij}:=\frac{\partial g_{i}}{\partial x_{j}}(x),\quad\forall\ x\in\mathbb{C}^{m}.$
Let $\hat{x}$ be a solution to the system of equations associated with the ideal generators. This solution has multiplicity greater than one, if and only if, the matrix $J(\hat{x})$ is rank deficient.
For “Complex 1”, $J(\hat{x},\hat{z})$ has size $30\times 51$ and rank 29. This implies that “Complex 1” has multiplicity greater than one, as we wanted to prove. Calculations are done with M2 code solMultiplicty.m2.
11.6 Cartesian coordinates on the sphere
@@
Equator yes 3 (×2)($\times$ 2)( × 2 ) yes 2
1:5 yes $\frac{6}{5}$ (×1)($\times$ 1)( × 1 ), $\frac{12}{5}$ (×2)($\times$ 2)( × 2 ) yes 3
1:4:1 yes 2 (×3)($\times$ 3)( × 3 ) yes 3
3:3 ($\sqrt{6}$) yes 2.28(×1$\simeq$ 2.28($\times$ 1)≃ 2.28 ( × 1 ), 1.86(×2$\simeq$ 1.86($\times$ 2)≃ 1.86 ( × 2 ) yes 3
3:3 ($-\sqrt{6}$) yes 9.48(×1$\simeq$-9.48($\times$ 1)≃ - 9.48 ( × 1 ), 7.74(×2$\simeq$ 7.74($\times$ 2)≃ 7.74 ( × 2 ) no 3
@@
Real 1 yes $\frac{4}{3}(\times 3)$, 2(×1)2($\times$ 1)2 ( × 1 ) yes 4
@@
11.7 Energies and solution in $S^{3}$ for six points
@@
Complex 2 non real coord. 3 180
@@
For each sphere $S^{d-1}$, the optimal configuration $X$ has the maximum number of conjugate symmetries (minimum orbit size). This is also true for $n=4$ and $n=5$, as can be seen in Tables 4 and 5. Also, within each sphere $S^{d-1}$, the product energy increases as the conjugate symmetries increase (decreases with the orbit size); the only exception being Real 2 and Real 3.
12 Geometric interpretation on $S^{3}$
Real 4, which is the solution on $S^{3}\subset\mathbb{R}^{4}$ for six points, consists of two equilateral triangles, each inscribed in a copy of $S^{1}$ lying in orthogonal spaces. In particular, from the cartesian coordinates given in (10), we see that points $w_{0}$, $w_{1}$ and $w_{2}$ form an equilateral triangle, inscribed in an $S^{1}$. This is because they have the same pairwise angles: $\cos^{-1}(-1/2)=\frac{2\pi}{3}$, and belong to the same plane through the origin: $w_{0}+w_{1}=-w_{2}$. The same happens with points $w_{3}$, $w_{4}$ and $w_{5}$, which form an equilateral triangle, inscribed in another $S^{1}$. Finally, these circumferences lie in orthogonal spaces:
@@ 12 Geometric interpretation on $S^{3}$
For the other critical configurations, they can be viewed as given by the fibration of $S^{3}$ by spheres. Real 1 is, in fact, analogous to the 1:4:1 configuration of $S^{2}$: it has the poles and the optimal configuration for 4 points in the equatorial sphere. Real 3 is, in turn, analogous to the 1:5 configuration of $S^{2}$: it has a pole and the optimal configuration for 5 points in the corresponding sphere. Real 2 has no analogous configuration on $S^{2}$. It has 4 points on the Equator of a sphere and the optimal configuration for 2 points in another sphere, with the peculiarity that the line that passes through these two points is orthogonal to the plane of the 4 point Equator.
13 Classification of critical configurations
@@
$\frac{\partial L}{\partial w_{k}}=-\sum_{j=1,j\neq k}^{n}\frac{2\left(w_{k}-w_{j}\right)}{\|w_{k}-w_{j}\|^{2}}+(n-1)w_{k},\quad\forall\ k=1,\ldots,n.$
@@
$\frac{\partial L}{\partial w_{k}}=\sum_{j=1,j\neq k}^{n}f(w_{k}-w_{j})+(n-1)w_{k},\quad\forall\ k=1,\ldots,n.$
@@
$f(w):=-\frac{2w}{\|w\|^{2}}=-\frac{2(x,y,z)}{x^{2}+y^{2}+z^{2}},\quad w:=(x,y,z).$
@@
$(H_{L})_{ii}:=\frac{\partial L}{\partial w_{i}^{2}}=\sum_{j=1,j\neq i}J_{f}(w_{i}-w_{j})+(n-1)I_{3},\quad\forall\ i.$
$(H_{L})_{ij}:=\frac{\partial L}{\partial w_{i}w_{j}}=-J_{f}(w_{i}-w_{j}),\quad\forall\ i\neq j.$
@@
$h_{L}:=V^{T}H_{L}V,\quad h_{L}\in\mathbb{R}^{n(d-1)\times n(d-1)};$
@@
    if $h_{L}$ has at least one negative eigenvalue, $w$ is a saddle point.
@@
    if $h_{L}$ has eigenvalue zero with multiplicity exactly $d(d-1)/2$, and all other eigenvalues of $h_{L}$ are positive, then $w$ is a local minimum (possibly global).
@@
    If $w$ is a saddle point in $S^{k}$, it is also a saddle in $S^{l}$, $\forall\ l\geq k$.
@@
    As the energy remains constant under rotations, $h_{L}$ will always have zero as an eigenvalue, with multiplicity at least the dimension of the orthogonal group: $|O(d)|=d(d-1)/2$. When this inequality is exact, zero is associated to rotations only. If the other eigenvalues are positive, then the Hessian classifies and the critical configuration is a local minima.

### v3

@@ 4.2 Six points
$X_{1:5}=\begin{pmatrix}1&A&A&A&A&A\\ A&1&B&C&C&B\\ A&B&1&B&C&C\\ A&C&B&1&B&C\\ A&C&C&B&1&B\\ A&B&C&C&B&1\end{pmatrix},\qquad\begin{array}[]{l}A:=-\frac{1}{5},\\ B:=\frac{-5+6\sqrt{5}}{25},\\ C:=\frac{-5-6\sqrt{5}}{25}\end{array}.$
$X_{1:4:1}=\begin{pmatrix}1&A&0&0&0&0\\ A&1&0&0&0&0\\ 0&0&1&0&A&0\\ 0&0&0&1&0&A\\ 0&0&A&0&1&0\\ 0&0&0&A&0&1\end{pmatrix},\qquad A:=-1.$
$X_{\text{Eq.}}=\begin{pmatrix}1&A&B&C&B&A\\ A&1&A&B&C&B\\ B&A&1&A&B&C\\ C&B&A&1&A&B\\ B&C&B&A&1&A\\ A&B&C&B&A&1\end{pmatrix},\qquad\begin{array}[]{l}A:=\frac{1}{2},\\ B:=-\frac{1}{2},\\ C:=-1\end{array}$
@@
4.2.1 Configuration 3:3 and its real conjugate
@@
$X_{3:3}=\begin{pmatrix}1&A&A&B&C&C\\ A&1&A&C&B&C\\ A&A&1&C&C&B\\ B&C&C&1&A&A\\ C&B&C&A&1&A\\ C&C&B&A&A&1\end{pmatrix},\qquad\begin{array}[]{l}A:=\frac{-7+3\sqrt{6}}{5}\\ B:=\frac{11-4\sqrt{6}}{5}\\ C:=\frac{-1-\sqrt{6}}{5}\end{array}.$
To verify that $X_{3:3}$ is critical, we replace its values in the polynomial equations using M2. As the matrix depends on $\sqrt{6}$, we define a polynomial ring with coefficients in $k:=\mathbb{Q}[t]/(t^{2}-6)$. This introduces a variable $t$, such that $t^{2}=6$. We then replace $\sqrt{6}$ by $t$ in $X_{3:3}$, and check that its entries verify the equations in $k[x_{ij},z_{ij}]$. As this is true for every $t$ such that $t^{2}=6$, it implies that if we change $\sqrt{6}$ by $\sqrt{6}$ in $X_{3:3}$, we also obtain a solution to the polynomial equations. We call this the “real conjugate” of 3:3, with matrix:
$X_{\overline{3:3}}=\begin{pmatrix}1&\overline{A}&\overline{A}&\overline{B}&\overline{C}&\overline{C}\\ \overline{A}&1&\overline{A}&\overline{C}&\overline{B}&\overline{C}\\ \overline{A}&\overline{A}&1&\overline{C}&\overline{C}&\overline{B}\\ \overline{B}&\overline{C}&\overline{C}&1&\overline{A}&\overline{A}\\ \overline{C}&\overline{B}&\overline{C}&\overline{A}&1&\overline{A}\\ \overline{C}&\overline{C}&\overline{B}&\overline{A}&\overline{A}&1\end{pmatrix},\qquad\begin{array}[]{l}\overline{A}:=\frac{-7-3\sqrt{6}}{5}\simeq-2.87\\ \overline{B}:=\frac{11+4\sqrt{6}}{5}\simeq 4.16\\ \overline{C}:=\frac{-1+\sqrt{6}}{5}\simeq 0.29\end{array}.$
As $X_{\overline{3:3}}$ has entries with absolute value greater than one, it is not the dot product of points on a unit sphere (it corresponds to a complex critical configuration). Note also that we already encountered configurations with non-rational values: the Equator for $n=5$, and 1:5; both depending on $\sqrt{5}$. However, in those cases, changing $\sqrt{5}$ by $\sqrt{5}$ gives a matrix in the same orbit as the original.
@@
4.2.2 Possible values for the dot product
@@ 8 $(5x_{45}^{2}+1)^{2}$ $\pm\frac{i}{\sqrt{5}}$
9 $5x_{45}^{2}-22x_{45}+5$ $\frac{11\pm 4\sqrt{6}}{5}$
10 $5x_{45}^{2}+2x_{45}-1$ $\frac{-1\pm\sqrt{6}}{5}$
11 $5x_{45}^{2}+14x_{45}-1$ $\frac{-7\pm 3\sqrt{6}}{5}$
12 $25x_{45}^{2}+28x_{45}+19$ $\frac{-14\pm i3\sqrt{31}}{25}$
13 $125x_{45}^{2}+50x_{45}-31$ $\frac{-5\pm 6\sqrt{5}}{25}$
14 $100x_{45}^{4}+95x_{45}^{3}-21x_{45}^{2}-22x_{45}+10$ 4 complex
15 $250x_{45}^{4}+110x_{45}^{3}-21x_{45}^{2}-19x_{45}+4$ 4 complex
16 $400x_{45}^{4}+488x_{45}^{3}-111x_{45}^{2}-196x_{45}+67$ 4 complex
17 $3x_{45}+1$ -1/3
18 $5x_{45}+4$ -4/5
19 $0x_{45}-1 1/1$0
20 $25x_{45}-1$ 1/25
21 $25x_{45}+11$ -11/25
22 $25x_{45}+23$ -23/25
Note that finding a Gröbner basis with an elimination order is usually more expensive than using a grevlex order, both in execution time and memory usage. However, since our ideal is zero dimensional, the projection could be done with any Gröbner basis, by means of working in the finite dimensional vector space $R/I$ [14][Ch. 2, Exercise 2]. As an alternative, a change of monomial order could be applied efficiently by using the FGLM algorithm (for example from grevlex to lex) [18]. In particular, the same Gröbner basis used to count solutions could be used to find the possible values of the solutions coordinates, without too significant additional effort.
4.2.3 Minimal primes and new solutions
@@
4.2.4 Orbits including new solutions
@@
Difference 938848=90938-848=90
4.2.5 Multiplicity of “Complex 1”
Proposition 4 ([23, Theorem 5.1]).
Consider an ideal $I\subset\mathbb{Q}[x_{1},\ldots,x_{m}]$, generated by polynomials $\{g_{1},\ldots,g_{r}\}$. Define the Jacobian matrix of the ideal generators as:
$J(x)\in\mathbb{R}^{r\times m}\qquad/\qquad J(x)_{ij}:=\frac{\partial g_{i}}{\partial x_{j}}(x),\qquad\forall\ x\in\mathbb{C}^{m}.$
Let $\hat{x}$ be a solution to the system of equations associated with the ideal generators. This solution has multiplicity greater than one, if and only if, the matrix J(x^)J($\hat${x}) is rank deficient.
[…]
4.2.6 Cartesian coordinates on the sphere
@@
Equator yes 3 $(\times 2)$ yes 2
1:5 yes $\frac{6}{5}$ $(\times 1)$, $\frac{12}{5}$ $(\times 2)$ yes 3
1:4:1 yes 2 $(\times 3)$ yes 3
3:3 ($\sqrt{6}$) yes $\simeq 2.28(\times 1)$, $\simeq 1.86(\times 2)$ yes 3
3:3 ($\sqrt{6}$) yes $\simeq-9.48(\times 1)$, $\simeq 7.74(\times 2)$ no 3
@@
Real 1 yes $\frac{4}{3}(\times 3)$, $2(\times 1)$ yes 4
@@
4.2.7 Energies and solution in $S^{3}$ for six points
@@
Complex 2 non real coord. 3 360
@@
For each sphere $S^{d-1}$, the optimal configuration $X$ has the maximum number of conjugate symmetries (minimum orbit size). This is also true for $n=4$ and $n=5$, as can be seen in Tables 5 and 6. Also, within each sphere $S^{d-1}$, the product energy increases as the conjugate symmetries increase (decreases with the orbit size); the only exception being Real 2 and Real 3.
4.3 Geometric interpretation on $S^{3}$
Real 4, which is the solution on $S^{3}\subset\mathbb{R}^{4}$ for six points, consists of two equilateral triangles, each inscribed in a copy of $S^{1}$ lying in orthogonal spaces. In particular, from the cartesian coordinates given in Equation (11) of the appendix, we see that points $w_{0}$, $w_{1}$ and $w_{2}$ form an equilateral triangle, inscribed in an $S^{1}$. This is because they have the same pairwise angles: $\cos^{-1}(-1/2)=\frac{2\pi}{3}$, and belong to the same plane through the origin: $w_{0}+w_{1}=-w_{2}$. The same happens with points $w_{3}$, $w_{4}$ and $w_{5}$, which form an equilateral triangle, inscribed in another $S^{1}$. Finally, these circumferences lie in orthogonal spaces:
@@ 4.3 Geometric interpretation on $S^{3}$
The other critical configurations can be viewed as given by the fibration of $S^{3}$ by spheres, as follows. Real 1 is, in fact, analogous to the 1:4:1 configuration of $S^{2}$: it has the poles and the optimal configuration for 4 points in the equatorial sphere. Real 3 is, in turn, analogous to the 1:5 configuration of $S^{2}$: it has a pole and the optimal configuration for 5 points in the corresponding sphere. Real 2 has no analogous configuration on $S^{2}$. It has 4 points on the Equator of a sphere and the optimal configuration for 2 points in another sphere, with the peculiarity that the line that passes through these two points is orthogonal to the plane of the 4 point Equator.
4.4 Classification of critical configurations
@@
$\frac{\partial L}{\partial w_{k}}=-\sum_{j=1,j\neq k}^{n}\frac{2\left(w_{k}-w_{j}\right)}{\|w_{k}-w_{j}\|^{2}}+(n-1)w_{k},\qquad\forall\ k=1,\ldots,n.$
@@
$\frac{\partial L}{\partial w_{k}}=\sum_{j=1,j\neq k}^{n}f(w_{k}-w_{j})+(n-1)w_{k},\qquad\forall\ k=1,\ldots,n,$
@@
$f(w):=-\frac{2w}{\|w\|^{2}}=-\frac{2(x,y,z)}{x^{2}+y^{2}+z^{2}},\qquad w:=(x,y,z).$
@@
$(H_{L})_{ii}:=\frac{\partial L}{\partial w_{i}^{2}}=\sum_{j=1,j\neq i}J_{f}(w_{i}-w_{j})+(n-1)\textrm{Id}_{3},\qquad\forall\ i.$
$(H_{L})_{ij}:=\frac{\partial L}{\partial w_{i}w_{j}}=-J_{f}(w_{i}-w_{j}),\qquad\forall\ i\neq j.$
@@
$h_{L}:=V^{T}H_{L}V,\qquad h_{L}\in\mathbb{R}^{n(d-1)\times n(d-1)};$
@@
    if $h_{L}$ has at least one negative eigenvalue, then $w$ is a saddle configuration.
@@
    if $h_{L}$ has eigenvalue zero with multiplicity exactly d(d1)/2d(d-1)/2, and all other eigenvalues of $h_{L}$ are positive, then $w$ is a local minimum (possibly global).
@@
    If $w$ is a saddle configuration in $S^{k}$, it is also a saddle in $S^{l}$, $\forall\ l\geq k$.
@@
    As the energy remains constant under rotations, $h_{L}$ will always have zero as an eigenvalue, with multiplicity at least the dimension of the orthogonal group: |O(d)|=d(d1)/2|O(d)|=d(d-1)/2. When this inequality is exact, zero is associated to rotations only. If the other eigenvalues are positive, then the Hessian classifies and the critical configuration is a local minima.